OpenStudy (anonymous):

Perform the indicated operation. Simplify the answer when possible. square root 2 + 2 square root 98 - 4 square root 8

OpenStudy (mathmale):

You may find that "square root of 2" is a factor common to all three of your terms. Please note that 98 = 2*49, and that 49 is a perfect square. Similarly, 8=2*4, and 4 is a perfect square. See whether you can use these hints to factor the given expression. Note: I'm getting off the 'Net now for a while.

OpenStudy (anonymous):

I don't understand

jimthompson5910 (jim_thompson5910):

do you agree that 98 = 2*49 ?

OpenStudy (anonymous):

I did that but I got this